# Probabilistic approach to feedback control enhances multi-legged locomotion on rugged landscapes

### v1

@@ Abstract
Achieving robust legged locomotion on complex terrains poses challenges due to the high uncertainty in robot-environment interactions. Recent advances in bipedal and quadrupedal robots demonstrate good mobility on rugged terrains but rely heavily on sensors for stability due to low static stability from a high center of mass and a narrow base of support[1]. We hypothesize that a multi-legged robotic system can leverage morphological redundancy from additional legs to minimize sensing requirements when traversing challenging terrains. Studies suggest [2, 3] that a multi-legged system with sufficient legs can reliably navigate noisy landscapes without sensing and control, albeit at a low speed of up to 0.1 body lengths per cycle (BLC). However, the control framework to enhance speed on challenging terrains remains underexplored due to the complex environmental interactions, making it difficult to identify the key parameters to control in these high-degree-of-freedom systems. Here, we present a bio-inspired vertical body undulation wave as a novel approach to mitigate environmental disturbances affecting robot speed, supported by experiments and probabilistic models. Finally, we introduce a control framework which monitors foot-ground contact patterns on rugose landscapes using binary foot-ground contact sensors to estimate terrain rugosity. The controller adjusts the vertical body wave based on the deviation of the limb’s averaged actual-to-ideal foot-ground contact ratio, achieving a significant enhancement of up to 0.235 BLC on rugose laboratory terrain. We observed a $\sim$ 50% increase in speed and a $\sim$ 40% reduction in speed variance compared to the open-loop controller. Additionally, the controller operates in complex terrains outside the lab, including pine straw, robot-sized rocks, mud, and leaves.
@@ I Introduction
Legged robots offer a compelling alternative to wheeled robots, particularly when navigating unstructured terrain. Recent advances in few-legged robots demonstrate their adaptability to diverse and unstructured terrains. Bipedal robots excel at obstacle avoidance and stair climbing [4, 5, 6], but their static instability necessitates substantial effort to maintain balance in an upright posture. Disruptions in body or leg trajectories can lead to instability [7, 8], limiting mobility on highly uneven terrains. Quadruped robots, known for greater stability, perform well on challenging terrains like snow, wet moss, mud, and rocky surfaces [9, 10, 11]. However, they rely on advanced control algorithms and complex sensing systems, including force sensors and accelerometers, to analyze foot-environment interactions and maintain balance [12, 13, 14, 15]. Additionally, they utilize cameras and LiDAR for estimating terrain geometric features for foothold planning [16, 17, 18, 19, 20]. This adds complexity to the sensing system and limits the robot’s scalability in confined spaces. In contrast, hexapod robots, with two additional legs, exhibit more robust mobility on rugged terrain[21, 22]. Multi-legged robots with more than six legs exhibit excellent static stability due to a low center of mass and broad base of support [2, 3, 23], minimizing the effort required to maintain balance. This stability gives the system significant potential for traversing rugged environments, allowing for a much simpler sensory system compared to fewer-legged robots.
@@
To address these limitations and improve the effectiveness of multi-legged locomotion in complex environments, designing a feedback control framework is essential. Specifically, designing an simple feedback controller to enhance the robot’s performance with minimal sensing is a challenging yet fascinating problem. To the best of our knowledge, feedback control for such multi-legged robotic systems has been minimally exploration [24, 25], largely due to their high degrees of freedom (over 25), complex environmental interactions, and unclear key parameters requiring control.
@@
To summarize our contribution, we first introduce a robotic system capable of generating coordinated body undulation and leg-stepping waves to enable forward motion. Next, we develop a binary contact sensing system to detect foot-ground contact. We then propose two probabilistic models to quantify the benefits of vertical wave modulation for multi-legged locomotion on rugged terrains. The first model predicts the robot’s forward speed based on the actual-to-ideal foot-ground contact ratio, while the second model estimates this contact ratio given the terrain distribution and vertical wave amplitude. By combining these models, we demonstrate that: 1) a higher contact ratio generally leads to increased speeds, and 2) appropriately adjusting the vertical wave amplitude can effectively achieve this higher contact ratio. Utilizing these models, we design a feedback controller that allows the robot to optimize its vertical amplitude based on terrain roughness, estimated in real-time through contact ratio measurements. Finally, we validate the controller’s effectiveness through laboratory and outdoor experiments.
@@ II Robot hardware design
Our robotic system’s thrust primarily originates from foot-ground interaction, in contrast to body-driven propulsion, which is more typical in robots operating continuously in environments like fluids or granular media [36, 37, 38]. When a leg contacts the ground and slips backward, it generates Coulomb friction with a forward-directed force (Fig. 5.B). Terrain heterogeneity can disrupt the planned foot-ground contact, potentially reducing the robot’s forward speed. Therefore, tracking the contact state to mitigate perturbations with a controller is an effective approach to enhance the robot’s body forward velocity.
@@ II-A Robophysical Model
We incorporated compliant legs inspired by the passive morphology of centipedes observed during their interactions with its surroundings [35], enhancing the robots’ ability to traverse rough terrain. Each leg consists of a two-bar linkage with a unidirectional hinge (knee joint) that bends opposite to the direction of forward motion. An elastic rubber band (Alliance Rubber 26324 Advantage Rubber Bands Size 32) restores the linkage to a neutral position after the leg encounters obstacles, aligning the links. This design ensures sufficient thrust during retraction, which is crucial for navigating uneven surfaces. The directional compliance facilitates a more evenly distributed contact area [39], improving the robot’s ability to traverse obstacles without significantly disrupting its gait.
Each leg connects at a pivot joint, allowing it to pitch up to a maximum angle of $60^{\circ}$ (7 cm above the ground). A motor (AX-12A, DYNAMIXEL) cable-driven system alternates lifting opposing legs: rotating counterclockwise raises the left leg while the right remains grounded, and rotating clockwise raises the right leg while the left remains grounded. When cable tension decreases, another rubber band restores the leg to its original pitch angle. A second motor (AX-12A, DYNAMIXEL) controls the yaw ($\theta_{leg}$) of the leg assembly, facilitating leg swing. Lastly, a 2-DoF motor (2XL430-W250, DYNAMIXEL) controls the yaw ($\theta_{body}$, Fig. 2.C.1) and pitch ($\theta_{v}$, Fig. 2.C.2) of the module, and connecting multiple modules facilitates lateral and vertical undulation (Fig.2.C.1 and C.2).
@@ II-B Binary contact sensing system design
We implemented a low-bandwidth binary contact sensor system (Fig. 3.A) to monitor foot-ground interaction for each leg, allowing us to assess terrain heterogeneity and using this information as a feedback into our control system. Contact capacitive sensors (MPR121) embedded at the tip (highlighted in yellow in Fig.3.A) of each leg detect capacitance variance. The toe((highlighted in pink in Fig.3.A)) has a slight range of linear motion, resulting in minimal capacitance when the leg is suspended and maximal capacitance when it is grounded. We used raw capacitance levels process through a threshold detection algorithm to determine contact. Figure.3.C illustrates experimental and predicted contact states during the forward gait on level ground for an 8-leg robot, demonstrating strong alignment between the two.
@@ III-A Leg and body wave coordination
Consequently, the gaits of multi-legged locomotors, achieved through the superposition of body wave and leg waves, are characterized by the phases of contact, $\phi_{c}$, and lateral body undulation, $\tau_{b}$. As outlined in [23], the optimal body-leg coordination, which phases body undulation to facilitate leg retraction, is given by $\phi_{c}=\tau_{b}-(\xi/N+1/2)\pi$. We fixed both $\Theta_{body}$ and $\Theta_{leg}$ at $30^{\circ}$ for all the experiments, as this coordination effectively mitigate environmental disturbances that impacted the robot’s speed [2].
@@ IV-C Contact ratio $\gamma$
Figure 7: Vertical body undulation serves as a mechanism to alleviate foot-ground contact state distortion. A. Examples of contact loss on rough terrain. (left) Contact loss occurs when the leg’s length is insufficient to reach the ground. (right) Contact loss occurs when the flexible leg is deformed by shifts in terrain level. In this scenario, the supporting force $F_{r}$ diminishes as it transitions from supporting the body weight to generating torque on the rubber band. B. Vertical body undulation reduces the contact contamination. (left) By lowering the leg deeper, contact contamination is minimized. (right) Conversely, lifting the leg facilitates its return to its initial undeformed state, thereby reducing contact contamination. C. Sketch for deriving probabilistic models: An illustration depicting the geometric relationship between the robotic structure and variations in terrain height.
@@ IV-D Probabilistic model for correlation between robot speed and $\gamma$
Thus far, we have defined a parameter $\gamma$ that directly influences the thrust generation of the robot. As shown in Fig. 5.B, during the retraction period or stance phase, the robot’s leg makes contact with the ground and swings backward ($v_{\text{slip}}$), resulting in Coulomb friction components in the direction of forward speed ($\bar{v}$). The contact loss could decrease $\gamma$ and potentially reduce the forward speed of the robot. The subsequent objective is to elucidate the relationship between the velocity of the robot and $\gamma$. Our experimental results in Fig. 5.C demonstrate a positive correlation between the robot’s speed and $\gamma$. Next, we derive a probabilistic model to describe the correlation between the robot’s speed $\bar{v}$ and $\gamma$ in sequential order,
@@ IV-D1 Correlation between friction and $\gamma$
Figure 6.A.2 shows the probability distribution of the slipping angle $\phi$ for the first leg (Figure 4) of a 12-legged robot. To derive this distribution, we conducted three cycles of motion with the robot on flat ground, each repeated for 10 trials. The trajectory of the leg was meticulously tracked using our motion tracking system (OptiTrack). By computing the tangential direction of the nearest point on the trajectories, we obtained the distribution of slipping angles.
@@ IV-D2 Correlation between friction and $\bar{v}$
Here, $\delta\bar{F}$ represents the disturbance to the average friction, while $\delta{\bar{v}}$ denotes the disturbance to the speed. The constant $C$ is determined by the robot’s dimensions and its ideal steady-state speed $v_{open}$.
@@
$\frac{\bar{v}}{v_{open}}\approx 1.065\frac{\bar{F}^{{}^{\prime}}_{m}}{\mu F_{w}}$ (12)
@@ IV-D3 Correlation between $\bar{v}$ and $\gamma$
Figure 8: Theoretical predictions and experimental validation of probabilistic models for the $\gamma-A_{v}$ correlation. A. Definition of contact error: the probability of contact error, denoted as $P_{e}$, is the portion of contacts that deviate from their corresponding flat terrain contact states. B. Probabilistic predictions of $P_{e}$ and $\gamma$ for various vertical amplitudes ($A_{v}$) across terrains with different rugosity. C. Experimental validation of the probabilistic models.
@@ V-A1 Gait without vertical wave
If $\triangle H>0$, it indicates that the terrain at the next time step is higher than the current one. In this case, the situation is more complex because the leg can still make contact with the ground. Our robot’s flexible legs can bend when encountering higher terrain. We assume a contact error occurs if the robot leg deforms upon encountering higher terrain (see Fig. 7.B, right). The contact is then restored by the leg’s retraction movement, represented by the retraction distance $D_{s}(t)$. During the retraction process, we assume the leg remains fixed at the foot-ground contact point (see Fig. 7.C, right). Then, we discretize the retraction period over a motion cycle into $m$ time steps. For a given time step $t_{i}$ during the retraction period of the robot’s leg, we define the maximum height difference that can be recovered by retraction as $\delta h(t_{i})^{{}^{\prime}}$, which is computed as:
@@ V-A3 Contact ratio $\gamma$ prediction
As a reminder, the contact ratio, denoted as $\gamma$, is defined as the resemblance between the actual contact state and the ideal contact state during the protraction period. The computation of the contact ratio is expressed by the formula:

### v2

@@ Abstract
Achieving robust legged locomotion on complex terrains poses challenges due to the high uncertainty in robot-environment interactions. Recent advances in bipedal and quadrupedal robots demonstrate good mobility on rugged terrains but rely heavily on sensors for stability due to low static stability from a high center of mass and a narrow base of support[1]. We hypothesize that a multi-legged robotic system can leverage morphological redundancy from additional legs to minimize sensing requirements when traversing challenging terrains. Studies suggest [2, 3] that a multi-legged system with sufficient legs can reliably navigate noisy landscapes without sensing and control, albeit at a low speed of up to 0.1 body lengths per cycle (BLC). However, the feedback control framework to enhance speed of multi-legged robots on challenging terrains remains underexplored due to diverse environmental interactions. Such complexity makes it difficult to identify the key parameters to control in these high-degree-of-freedom systems. Here, in laboratory and field experiments we demonstrate that a vertical body undulation wave helps to mitigate environmental disturbances affecting robot speed; these observations are supported by a probabilistic model. Using such insights, we introduce a control framework which monitors foot-ground contact patterns on rugose landscapes using binary foot-ground contact sensors to estimate terrain rugosity. The controller adjusts the vertical body wave based on the deviation of the limb’s averaged actual-to-ideal foot-ground contact ratio, achieving a significant enhancement of up to 0.235 BLC on rugose laboratory terrain. We observed a 50% to 60% increase in speed and a 30% to 50% reduction in speed variance compared to the open-loop controller. Additionally, the controller operates in complex terrains outside the lab, including pine straw, robot-sized rocks, mud, and leaves.
@@ I Introduction
Legged robots offer a compelling alternative to wheeled robots, particularly when navigating unstructured terrain. Recent advances in few-legged robots demonstrate their adaptability to diverse and unstructured terrains. Bipedal robots excel at obstacle avoidance and stair climbing [4, 5, 6], but their static instability necessitates substantial effort to maintain balance in an upright posture. Disruptions in body or leg trajectories can lead to instability [7, 8], limiting mobility on highly uneven terrains. Quadruped robots, known for greater stability, perform well on challenging terrains like snow, wet moss, mud, and rocky surfaces [9, 10, 11]. However, many advanced robotic systems rely on force sensors, accelerometers [12, 13, 14, 15], cameras, and LiDAR [16, 17, 18, 19, 20] for analyzing foot-environment interactions and estimating terrain geometry. However, not all legged robots require these systems. Systems with high static stability and morphological redundancy, such as hexapods [21, 22] or robots with more legs [2, 3, 23], often traverse rugged terrain effectively with simpler sensing and control frameworks. These simpler systems minimize sensory complexity while still achieving reliable locomotion.
@@
To address these limitations and improve the effectiveness of multi-legged locomotion in complex environments, designing a feedback control framework is essential. Specifically, designing an simple feedback controller to enhance the robot’s performance with minimal sensing is a challenging yet fascinating problem. To the best of our knowledge, feedback control for such multi-legged robotic systems remains a challenging area of research, particularly due to their high degrees of freedom (over 25), complex environmental interactions, and the difficulty of identifying key parameters requiring control. While several studies [24, 25] have proposed different approaches, these efforts are often constrained by computational complexity, energy efficiency concerns, and limited adaptability to highly unstructured terrains.
@@
To summarize our contribution, we first introduce a robotic system capable of generating coordinated body undulation and leg-stepping waves to enable forward motion. Next, we develop a binary contact sensing system to detect foot-ground contact. We then propose two probabilistic models to quantify the benefits of vertical wave modulation for multi-legged locomotion on rugged terrains. The first model predicts the robot’s forward speed based on the actual-to-ideal foot-ground contact ratio, while the second model estimates this contact ratio given the terrain distribution and vertical wave amplitude. By combining these models, we demonstrate that: 1) a higher contact ratio generally leads to increased speeds, and 2) appropriately adjusting the vertical wave amplitude can effectively achieve this higher contact ratio. Utilizing these models, we design a feedback controller that allows the robot to optimize its vertical amplitude based on terrain roughness, estimated in real-time through contact ratio measurements. While our models and experiments primarily focus on terrains with varying height differences — one of the key causes of contact loss — the proposed framework is designed to address a broader class of rugged landscapes where such vertical irregularities are prevalent. We discuss the generalizability and limitations of this approach in Section VIII. Finally, we validate the controller’s effectiveness through laboratory and outdoor experiments.
@@ II Robot hardware design
Our robotic system’s thrust primarily originates from foot-ground interaction, in contrast to body-driven propulsion, which is more typical in robots operating continuously in environments like fluids or granular media [36, 37, 38]. When a leg contacts the ground and slips backward, it generates Coulomb friction with a forward-directed force. Terrain heterogeneity can disrupt the planned foot-ground contact, potentially reducing the robot’s forward speed. Therefore, tracking the contact state to mitigate perturbations with a controller is an effective approach to enhance the robot’s body forward velocity.
@@ II-A Robophysical Model
We incorporated compliant legs inspired by the passive morphology of centipedes observed during their interactions with its surroundings [35], enhancing the robots’ ability to traverse rough terrain. Each leg consists of a two-bar linkage with a unidirectional hinge (knee joint) that bends opposite to the direction of forward motion. An elastic rubber band (Alliance Rubber 26324 Advantage Rubber Bands Size 32) restores the linkage to a neutral position after the leg encounters obstacles, aligning the links. This design ensures sufficient thrust during retraction, which is crucial for navigating uneven surfaces. The directional compliance facilitates a more evenly distributed contact area [39], improving the robot’s ability to traverse obstacles without significantly disrupting its gait. Experiments comparing the performance of rigid and compliant legs on rough terrain can be found in the Appendix -D.
Each leg connects at a pivot joint, allowing it to pitch up to a maximum angle of $60^{\circ}$ (7 cm above the ground). A motor (AX-12A, DYNAMIXEL) cable-driven system alternates lifting opposing legs: rotating counterclockwise raises the left leg while the right remains grounded, and rotating clockwise raises the right leg while the left remains grounded. When cable tension decreases, another rubber band restores the leg to its original pitch angle. A second motor (AX-12A, DYNAMIXEL) controls the yaw ($\theta_{leg}$) of the leg assembly, facilitating leg swing. The yaw angle ($\theta_{leg}$) is defined as the shoulder angle for corresponding segment. Lastly, a 2-DoF motor (2XL430-W250, DYNAMIXEL) controls the yaw ($\theta_{body}$, Fig. 2.C.1) and pitch ($\theta_{v}$, Fig. 2.C.2) of the module, and connecting multiple modules facilitates lateral and vertical undulation (Fig.2.C.1 and C.2).
@@ II-B Binary contact sensing system design
We implemented a low-bandwidth binary contact sensor system (Fig. 3.A) to monitor foot-ground interaction for each leg, allowing us to assess terrain heterogeneity and using this information as a feedback into our control system. Contact capacitive sensors (MPR121) embedded at the tip (highlighted in yellow in Fig.3.A) of each leg detect capacitance variance. The toe((highlighted in pink in Fig.3.A)) has a slight range of linear motion, resulting in minimal capacitance when the leg is suspended and maximal capacitance when it is grounded. The analog value shows a significant difference between the suspended state (greater than 200) and the grounded state (less than 5). Therefore, we classify any analog value below 50 as indicating contact. Figure.3.C illustrates experimental and predicted contact states during the forward gait on level ground for an 8-leg robot, demonstrating strong alignment between the two.
@@ III-A Leg and body wave coordination
Consequently, the gaits of multi-legged locomotors, achieved through the superposition of body wave and leg waves, are characterized by the phases of contact, $\tau_{c}$, and lateral body undulation, $\tau_{b}$. As outlined in [23], the optimal body-leg coordination, which phases body undulation to facilitate leg retraction, is given by $\tau_{c}$$=\tau_{b}-(\xi/N+1/2)\pi$. We fixed both $\Theta_{body}$ and $\Theta_{leg}$ at $30^{\circ}$ for all the experiments, as this coordination effectively mitigate environmental disturbances that impacted the robot’s speed [2].
@@ IV-D Probabilistic model for correlation between robot speed and $\gamma$
Thus far, we have defined a parameter $\gamma$ that directly influences the thrust generation of the robot. As shown in Fig. 5.B, during the retraction period or stance phase, the robot’s leg makes contact with the ground and swings backward ($v_{\text{tip}}$), resulting in Coulomb friction components in the direction of forward speed ($\bar{v}$). The contact loss could decrease $\gamma$ and potentially reduce the forward speed of the robot. The subsequent objective is to elucidate the relationship between the velocity of the robot and $\gamma$. Our experimental results in Fig. 5.C demonstrate a positive correlation between the robot’s speed and $\gamma$. Next, we derive a probabilistic model to describe the correlation between the robot’s speed $\bar{v}$ and $\gamma$ in sequential order,
@@ IV-D1 Correlation between friction and $\gamma$
Figure 6.A.2 shows the probability distribution of the slipping angle $\phi$ for the first leg (Figure 4) of a 12-legged robot. To derive this distribution, we conducted three cycles of motion with the robot on flat ground, each repeated for 10 trials. The trajectory of the leg was meticulously tracked using our motion tracking system (OptiTrack; see Appendix). By computing the tangential direction of the nearest point on the trajectories, we obtained the distribution of slipping angles.
@@ IV-D2 Correlation between friction and $\bar{v}$
Here, $\delta\bar{F}$ represents the disturbance to the average friction, while $\delta{\bar{v}}$ $=\bar{v}-v_{open}$ denotes the disturbance to the speed. The constant $C$ is determined by the robot’s dimensions and its ideal steady-state speed $v_{open}$. Specifically, $C=\int^{\pi}_{0}\mu F_{w}/v_{open}\sin(\tan^{-1}(v_{open}/v_{x}(\tau))d\tau$. Here, $v_{x}(\tau)$ is the lateral speed of the leg tip over time and can be obtained by analyzing robot’s geometry [3].
@@
$\frac{\bar{v}}{v_{open}}\approx 1.065(\frac{\bar{F}^{{}^{\prime}}_{m}}{\mu F_{w}}+1)$ (12)
@@ V Vertical body motion mitigates environmental disturbance
Figure 7: Vertical body undulation serves as a mechanism to alleviate foot-ground contact state distortion. A. Examples of contact loss on rough terrain. (left) Contact loss occurs when the leg’s length is insufficient to reach the ground. (right) Contact loss occurs when the flexible leg is deformed by shifts in terrain level. In this scenario, the supporting force $F_{r}$ diminishes as it transitions from supporting the body weight to generating torque on the rubber band. B. Vertical body undulation reduces the contact contamination. (left) By lowering the leg deeper, contact contamination is minimized. (right) Conversely, lifting the leg facilitates its return to its initial undeformed state, thereby reducing contact contamination. C. Sketch for deriving probabilistic models: An illustration depicting the geometric relationship between the robotic structure and variations in terrain height. (left) This sketch illustrates the maximum depth ($h_{r}$) that the leg can reach with the vertical joint pitch motion. (right) This sketch demonstrates how leg retraction restores the leg from deformation to its straight state. Given the retraction distance $D_{s}$, the compliance (knee) joint is lifted by $\delta h^{\prime}$. Circles without fill on the dashed line represent the pivot joint for leg lifting, while solid circles indicate the knee joint for leg compliance. $h_{l,1}$ is distance between the pivot joint and knee joint while $h_{l,2}$ is the distance between the knee joint and leg tip. Squares represent the stroke for leg retraction, showing the range of leg swing from anterior to posterior during the stance phase of one motion cycle.
@@ V-A1 Gait without vertical wave
If $\triangle H>0$, it indicates that the terrain at the next time step is higher than the current one. In this case, the situation is more complex because the leg can still make contact with the ground. Our robot’s flexible legs can bend when encountering higher terrain. We assume a contact loss occurs if the robot leg deforms upon encountering higher terrain (see Fig. 7.B, right).
The contact is then restored by the leg’s retraction movement, represented by the retraction distance $D_{s}(t)$. During the retraction process, we assume the leg remains fixed at the foot-ground contact point (see Fig. 7.C, right). Then, we discretize the retraction period over a motion cycle into $m$ time steps. For a given time step $t_{i}$ during the retraction period of the robot’s leg, we define the maximum height difference that can be recovered by retraction as $\delta h(t_{i})^{{}^{\prime}}$, which is computed as:
@@
Note that in Section IV, Fig. 5A, we define contact loss as regions where contact is lost—specifically, transitions from contact (on flat ground) to no contact (on rugged terrain)—relative to the ideal contact pattern. In Section V, we extend the definition of contact loss to include both:
  • Contact $\rightarrow$ No contact: This results in a missed step, as shown in Fig. 7 (left).
  • Contact $\rightarrow$ Contact, but with leg deformation: Although the foot maintains contact with the ground, the leg fails to maintain its intended shape or posture. In the stance phase, where ground contact is planned, such deformation is considered a form of contact loss, as illustrated in Fig. 7 (right).
@@ V-A3 Contact ratio $\gamma$ prediction
As a reminder, the contact ratio, denoted as $\gamma$, is defined as the likelihood between the actual contact state and the ideal contact state during the retraction period. The computation of the contact ratio is expressed by the formula:
@@ V-A4 Experiment validation
Figure 8: Theoretical predictions and experimental validation of probabilistic models for the $\gamma-A_{v}$ correlation. A. Definition of contact error: the probability of contact error, denoted as $P_{e}$, is the portion of contacts that deviate from their corresponding flat terrain contact states. B. Probabilistic predictions of $P_{e}$ and $\gamma$ for various vertical amplitudes ($A_{v}$) across terrains with different rugosity. C. Experimental validation of the probabilistic models. We conducted 10 experimental trials for each parameters. The error bars represent the standard deviation.